Given p= 2.11 determine the pKa for 3-nitophenol and compare this value to the literature value
Answer #1

pKa = pH at which half the acid is in ionized form

so at pH 7.15, 50% of p-nitrophenol is ionized

use the Henderson-Hasselbalch equation:
pH = pKa + log10 ([A-]/[HA])
7.5 = 7.15 + log ([A-]/[HA])
.35 = log [A-]/[HA]
10^.35 = [A-]/[HA]
2.24 =[A-]/[HA]
So at pH 7.5, p-nitrophenoxide is present at an amount 2.24 times that of p-nitrophenol

2.24x + x = 100%
3.24x = 100
x = 30.9 = percent in protonated form
2.24x = 69.1 = percent in ionized form at pH 7.5